# LLM2Vec: Large Language Models Are Secretly Powerful Text Encoders

### v1

@@ 4.1 LLM2Vec helps models to capture information from future tokens
To analyze the extent to which LLM2Vec-transformed models incorporate information from future tokens, we adopt the analysis of Springer et al. (2024) and test how good models are at judging the similarity of sentences that have different meanings but share the same prefix.
@@
We evaluate on a synthetic dataset collected by Springer et al. (2024), which consists of 35 sentence triples $\{(q_{i}$, $s^{+}_{i}$, $s^{-}_{i})\}_{i=1}^{35}$ with $q_{i}=(A_{i},B_{i})$, $s^{+}_{i}=(A_{i},C_{i})$, and $s^{-}_{i}=(A_{i},D_{i})$, where $B_{i}$ and $C_{i}$ have a similar meaning to each other but $B_{i}$ and $D_{i}$ don’t. We compute a sequence representation for each of these sentences by pooling only over the first part of the sentence, i.e., $A_{i}$. We then compute the cosine similarity between the resulting embeddings. A model that incorporates information from future tokens ($B_{i}$, $C_{i}$, or $D_{i}$) in the representations of the prefix $A_{i}$ should assign a higher similarity to the positive example.
Results
@@
Based on these findings (we replicate these results for other inputs and other Mistral models in Appendix F) and the strong unsupervised results for Mistral-7B with bidirectional attention, we speculate that Mistral models are pre-trained with some form bidirectional attention, e.g., prefix language modeling (Raffel et al., 2020) – at least for some parts of its training. We leave a more detailed investigation of this intriguing behavior for future work.
@@ 5.1 LLM2Vec leads to strong performance on the MTEB leaderboard
For supervised training, we train on a replication of the public portion of the E5 dataset (Wang et al., 2023) curated by Springer et al. (2024). The dataset consists of approximately 1.5M samples and we provide details on its compilation in Section G.1. We follow standard practice and train the models with contrastive learning using hard negatives and in-batch negatives. We use LoRA fine-tuning for the supervised setting as well. The MNTP LoRA weights are merged into the base model, and the trainable LoRA weights are initialized with SimCSE weights. For LLM2Vec models that use just MNTP, the LoRA weights are randomly initialized. The training is performed for 1000 steps with a batch size of 512. We detail other hyperparameters in Section G.2.
@@ Results
Table 2: Supervised results on full MTEB benchmark. The best performing LLM2Vec model achieves a new SOTA performance among models trained only on publicly available data.
Figure 6: Results on the 15-task subset of MTEB during training. For all three models, applying LLM2Vec before supervised training leads to better performance with fewer steps.
Table 2 shows the results of our evaluation. For all three models, transforming a model with LLM2Vec leads to improved performance over the strong Uni + weighted mean baseline. Performing unsupervised SimCSE is less crucial for supervised training, and even leads to slightly worse performance for LLaMA-2-7B and Mistral-7B compared to just performing the MNTP step of LLM2Vec. However, as we will show in Section 5.2, LLM2Vec with MNTP and SimCSE is much more sample-efficient, and therefore crucial in compute- or data-constrained settings. Notably, our best model, Mistral-7B (LLM2Vec w/o SimCSE) leads to a new state-of-the-art performance among models trained only on publicly available data.
@@ 5.2 LLM2Vec leads to more sample-efficient training
To demonstrate the sample-efficiency of LLM2Vec-transformed models, we save a checkpoint every 25 training steps and evaluate them on our 15-task subset of MTEB.
@@ Results
As shown in Figure 6, LLM2Vec-transformed models achieve better performance earlier in training. This observation is consistent across all three models. For S-LLaMA-1.3B, the smallest of our three models, even performing just MNTP leads to considerably improved sample efficiency. These results are particularly encouraging for settings where it is hard to acquire high quality labeled data, a setting which we leave for future work.
@@ 6 Related Work
Initially, supervised methods primarily relied on tasks such as natural language inference or sentence similarity to train BERT-like models for producing sentence embeddings (Conneau et al., 2017; Reimers & Gurevych, 2019). Subsequently, BERT-like models have also been adapted to tasks like retrieval (Karpukhin et al., 2020; Khattab & Zaharia, 2020). More recent methods have further improved these models through a complex multi-stage learning pipeline that consists of large-scale weakly supervised contrastive training followed by multi-task fine-tuning (Ni et al., 2022; Wang et al., 2022a; Li et al., 2023; Xiao et al., 2023) Recent approaches have focused on enhancing the generalization and transferability of text embeddings using instructions (Su et al., 2023; Asai et al., 2023).
@@
While decoder-only LLMs have outperformed bidirectional encoders across a large variety of language understanding tasks (Brown et al., 2020; Touvron et al., 2023; Jiang et al., 2023, inter alia), their impact on sentence representation learning remains limited. The most common approaches in the literature use the final hidden state of the last token as the sentence embedding (Neelakantan et al., 2022; Ma et al., 2023; Wang et al., 2023). Very recent (and concurrent) works address the inherent limitation of LLMs for sentence representation – their causal attention mechanism.
Muennighoff et al. (2024) perform multi-task full fine-tuning using a combination of self-supervised language modeling with causal attention and supervised contrastive learning with bidirectional attention. In contrast, our proposed approach is much more computationally efficient, as it requires only parameter-efficient fine-tuning and 1000 gradient steps. Closest to our work is the concurrent work of Springer et al. (2024). They propose copying the input sequence and appending it to itself, which addresses the contextualization issue of causal attention as tokens in the copy of the input can now attend to “future” tokens in the previous sequence. While this performs well in practice, it significantly increases the computational cost at inference time, which can be particularly problematic for encoding longer documents. Our approach outperforms Springer et al. (2024), without inducing any additional computational overhead at inference time.
@@ 7 Conclusion
We present LLM2Vec, a strong unsupervised approach to transform any decoder-only LLMs into a (universal) text embedder. We perform an extensive evaluation on word- and sequence-level tasks and demonstrate the effectiveness of LLM2Vec in both unsupervised and supervised settings. Applying LLM2Vec to Mistral-7B achieves a new state-of-the-art performance on MTEB among unsupervised approaches as well as among supervised approaches that train only on publicly available data. Beyond our strong empirical contributions, we provide an extensive analysis of how LLM2Vec impacts the underlying model and reveal an intriguing property of Mistral-7B, which explains its strong out of-the-box performance with bidirectional attention. The simplicity of our approach, as well as its compute and sample efficiency, makes LLM2vec a promising solution for low-resource and compute-constrained scenarios and opens up several interesting avenues for future work.
@@ Appendix A Limitations
The advantages of small bidirectional encoders come at the cost of complex and computationally intensive training regimes (Li et al., 2023; Xiao et al., 2023; Li et al., 2023). In contrast, decoder-only models are much more sample-efficient and do not require large-scale contrastive pre-training (Wang et al., 2023). Moreover, the instruction following capabilities of decoder-only models make them strong contenders for building text embedding models that generalize to a wide range of tasks and domains without the need for expensive adaptation.
@@ D.1.1 MNTP training details
For our models, we perform a hyperparameter search to select the percentage of the masked tokens in a sequence choosing from $20\%,40\%,60\%,80\%$, and $90\%$. For each model, we take the best setup (i.e., masking probability and BERT vs. RoBERTa approach) based on the performance on SICK-R (Agirre et al., 2014) task from the MTEB dataset. This results in the following choices: for S-LLaMA-1.3B and LLaMA-2-7B, we apply BERT’s masking strategy with masking probability of $20\%$. For Mistral-7B, we apply RoBERTa’s masking strategy with probability of $80\%$.
We train all the models for 1000 steps with LoRA $r=16$ and $\alpha=32$, and we follow the same training parameters as RoBERTa MNTP training. When training large 7B models, we apply brain floating point (bfloat16) quantization, as well as FlashAttention-2 (Dao, 2024) and gradient checkpointing.
@@ D.1.2 SimCSE training details
Similar to MNTP, we train all models with LoRA $r=16$ and $\alpha=32$ for 1000 steps. For LLaMA-2-7B and Mistral-7B, we train with a batch size of 128. For S-LLaMA-1.3B, we use a batch size of 32. Additionally, when training LLaMA-2-7B and Mistral-7B, we apply brain floating point (bfloat16) quantization, FlashAttention-2 (Dao, 2024), and gradient checkpointing.
@@ D.1.3 Word-level training details
We use data from CoNLL-2003, consisting of roughly 14,000 training, 3,250 validation, and 3,450 test samples (Tjong Kim Sang & De Meulder, 2003). We train the classifier for 1,500 steps with a learning rate of $0.0005$ and batch size 8.
As the models we experiment with have sub-token based vocabularies, we calculate the embedding of a word by averaging the representations of all its sub-tokens. For example, for a sentence “$w_{1}\;w_{2}\;w_{3}$” which is tokenized as “$\text{BOS}\;\;t_{11}t_{12}\;\;t_{21}t_{22}t_{23}\;\;t_{31}$”, the representation of $w_{1}$, $w_{2}$, and $w_{3}$ will be computed as
@@ Appendix F More analysis results
$q$: the query
$s^{+}$: the positive sample
$s^{-}$: the negative sample
She loves to travel in summer, especially to cold destinations, avoiding hot and crowded places.
She loves to travel in summer, specifically to chilly locations, steering clear of warm, populous areas.
She loves to travel in summer, but prefers to visit hot and bustling tourist spots.
The cat often sits by the window, dreaming of chasing birds and enjoying the warm sunshine.
The cat often sits by the window, imagining bird pursuits and basking in the sunlight.
The cat often sits by the window, but is too lazy to dream of chasing anything.
He reads books every night, finding solace in fiction and escaping from the stresses of daily life.
He reads books every night, seeking comfort in stories and evading everyday tensions.
He reads books every night, yet he feels that non-fiction is more engaging and informative.
She paints landscapes on weekends, expressing her creativity through vibrant colors and abstract forms.
She paints landscapes on weekends, showcasing her artistic flair with lively hues and unconventional shapes.
She paints landscapes on weekends, preferring realistic and detailed depictions of nature.
@@
Figure 7 provide cosine similarity results for all three models. In addition to our LLM2Vec-transformed models, we also provide results for Echo embeddings.
@@ F.1 Additional plots
Figure 8 provides additional plots for the representation analysis using two different Wikipedia paragraphs. The trends closely follow those reported in Section 4. Figure 9 shows that the same behavior we observe for Mistral-7B-instruct-v0.2 also holds true for other variants of the Mistral-7B model. We take this as additional evidence that the Mistral-7B base model was trained with some form of bidirectional attention.
@@ G.1 E5 dataset
NLI
Given a premise, retrieve a hypothesis that is entailed by the premise
@@
DuReader
Given a Chinese search query, retrieve web passages that answer the question
ELI5
Provided a user question, retrieve the highest voted answers on Reddit ELI5 forum
FEVER
Given a claim, retrieve documents that support or refute the claim
HotpotQA
Given a multi-hop question, retrieve documents that can help answer the question
MIRACL
Given a question, retrieve Wikipedia passages that answer the question
MrTyDi
Given a question, retrieve Wikipedia passages that answer the question
MSMARCO Passage
Given a web search query, retrieve relevant passages that answer the query
MSMARCO Document
Given a web search query, retrieve relevant documents that answer the query
NQ
Given a question, retrieve Wikipedia passages that answer the question
QuoraDuplicates
Given a question, retrieve questions that are semantically equivalent to the given question
@@
SQuAD
Retrieve Wikipedia passages that answer the question
T2Ranking
Given a Chinese search query, retrieve web passages that answer the question
TriviaQA
Retrieve Wikipedia passages that answer the question
@@ G.2 Training details
All models are trained with LoRA $r=16$ and $\alpha=32$, brain floating point (bfloat16) quantization, gradient checkpointing, and FlashAttention-2 (Dao, 2024) to optimize GPU memory consumption. We train on 8 NVIDIA A100 GPUs with an effective batch size of 512 for 1000 steps using a maximum sequence length of 512 tokens. We use the Adam optimizer with a learning rate of $0.0002$ and a linear learning rate warm-up for the first 300 steps.

### v2

@@ 4.1 LLM2Vec helps models to capture information from future tokens
To analyze the extent to which LLM2Vec-transformed models incorporate information from future tokens, we adopt the analysis of Springer et al. (2024) and test how well the model performs at judging the similarity between sentences that share the same prefix.
@@
We evaluate on a synthetic dataset collected by Springer et al. (2024), which consists of 35 sentence triples $\{(q_{i}$, $s^{+}_{i}$, $s^{-}_{i})\}_{i=1}^{35}$ with $q_{i}=(A_{i},B_{i})$, $s^{+}_{i}=(A_{i},C_{i})$, and $s^{-}_{i}=(A_{i},D_{i})$, where $B_{i}$ and $C_{i}$ have a similar meaning but $B_{i}$ and $D_{i}$ don’t. We compute a sequence representation for each of these sentences by pooling only over the first part of the sentence, i.e., $A_{i}$. We then compute the cosine similarity between the resulting embeddings. A model that incorporates information from future tokens ($B_{i}$, $C_{i}$, or $D_{i}$) in the representations of the prefix $A_{i}$ should assign a higher similarity to the positive example.
Results
@@
Based on these findings (we replicate these results for other inputs and other variants of Mistral in Appendix F) and the strong unsupervised results for Mistral-7B with bidirectional attention, we speculate that Mistral models are pre-trained with some form bidirectional attention, e.g., prefix language modeling (Raffel et al., 2020) – at least for some parts of its training. We leave a more detailed investigation of this intriguing behavior for future work.
@@ 5.1 LLM2Vec leads to strong performance on the MTEB leaderboard
For supervised training, we train on a replication of the public portion of the E5 dataset (Wang et al., 2023) curated by Springer et al. (2024). The dataset consists of approximately 1.5M samples and we provide details on its compilation in Section G.1. We follow standard practice and train the models with contrastive learning using hard negatives and in-batch negatives. We use LoRA fine-tuning for supervised setting as well. The MNTP LoRA weights are merged into the base model, and the trainable LoRA weights are initialized with SimCSE weights. For LLM2Vec models that use just MNTP, the LoRA weights are randomly initialized. The training is performed for 1000 steps with a batch size of 512. We detail other hyperparameters in Section G.2.
@@ Results
Meta-LLaMA-3-8B
Uni + w. Mean 55.42 58.60 43.19 86.29 75.56 83.95 30.59 63.87
LLM2Vec (w/o SimCSE) 56.63 59.68 46.45 87.80 75.92 83.58 30.94 65.01
LLM2Vec 56.71 59.02 45.86 87.95 76.67 82.98 29.67 64.90
Table 2: Supervised results on full MTEB benchmark. The best performing LLM2Vec model Meta-LLaMA-3-8B + LLM2Vec (w/o SimCSE) achieves a new SOTA performance among models trained only on publicly available data.
(a) S-LLaMA-1.3B
(b) Llama-2-7B
(c) Mistral-7B
Figure 7: Results on the 15 task subset of MTEB during training of S-LLaMA-1.3B, LLaMA-2-7B, and Mistral-7B. For all three models, applying LLM2Vec before supervised training leads to better performance with less steps.
Table 2 shows the results of our evaluation. For all models, transforming a model with LLM2Vec leads to improved performance over the strong Uni + weighted mean baseline. As expected, performing unsupervised SimCSE is less crucial for supervised training, and even leads to slightly worse performance for LLaMA-2-7B, Mistral-7B, and Meta-LLaMA-3-8B compared to just performing the MNTP step of LLM2Vec (LLM2Vec w/o SimCSE). However, as we will show in Section 5.2, LLM2Vec with MNTP and SimCSE is much more sample-efficient, and therefore crucial in compute or data-constrained settings. Notably, our best model, Meta-LLaMA-3-8B + LLM2Vec (w/o SimCSE) leads to a new state-of-the-art performance among models trained only on publicly available data.
@@ 5.2 LLM2Vec leads to more sample-efficient training
To demonstrate the sample-efficiency of LLM2Vec-transformed models, we save a checkpoint every 25 training steps and evaluate them on our 15 task subset of MTEB.
@@ Results
As shown in Figure 7, LLM2Vec-transformed models reach better performance earlier in training. This observation is consistent across all three models. For S-LLaMA-1.3B, the smallest of our three models, even performing just MNTP leads to a considerably improved sample-efficiency. These results are particularly encouraging for settings where it is hard to acquire high quality labeled data, a setting which we leave for future work.
@@ 6 Related Work
Initially, supervised methods primarily relied on tasks such as natural language inference or sentence similarity to train BERT-like models for producing sentence embeddings (Conneau et al., 2017; Reimers & Gurevych, 2019). Subsequently, BERT-like models have also been adapted to tasks like retrieval (Karpukhin et al., 2020; Khattab & Zaharia, 2020). More recent methods have further improved these representations through a complex multi-stage learning pipeline that consists of large-scale weakly supervised contrastive training followed by multi-task fine-tuning (Ni et al., 2022; Wang et al., 2022a; Li et al., 2023a; Xiao et al., 2023) Recent approaches have focused on enhancing the generalization and transferability of text embeddings using instructions (Su et al., 2023; Asai et al., 2023).
@@
While decoder-only LLMs have outperformed bidirectional encoders across a large variety of language understanding tasks (Brown et al., 2020; Touvron et al., 2023; Jiang et al., 2023a, inter alia), their impact on sentence representation learning remains limited. The most common approaches in literature use the final hidden state of the last token as the sentence embedding (Neelakantan et al., 2022; Ma et al., 2023; Wang et al., 2023).
There are few works that explore the limitations of using a causal attention mask when adapting decoder-only LLMs for text classification and sentence representation tasks. Li et al. (2023b) experiment with removing the causal mask of Llama-2 during supervised fine-tuning for text classification and NER tasks. Similarly, Dukić & Šnajder (2024) enable bidirectional attention for a group of layers during supervised fine-tuning on NER and chunking. In the context of sentence representation learning, Li & Li (2024) explore enabling bidirectional attention in the last layer of a decoder-only model during supervised contrastive fine-tuning on STS tasks.
Concurrent to our work, several works have focused on converting decoder-only-LLMs to text encoders in supervised and unsupervised manner. Jiang et al. (2023b) and Lei et al. (2024) prompt the language model to summarize the input text in one word, and take the last layer’s hidden embedding for the last token as the text’s representation. Muennighoff et al. (2024) perform multi-task full fine-tuning using a combination of self-supervised language modeling with causal attention and supervised contrastive learning with bidirectional attention. In contrast, our proposed approach is much more computationally efficient, as it requires only parameter-efficient fine-tuning and 1000 gradient steps. Closest to our work is the concurrent work of Springer et al. (2024). They propose to copy the input sequence and append it to itself, which addresses the contextualization issue of causal attention as tokens in the copy of the input can now attend to ”future” tokens in the previous sequence. While this performs well in practice, it significantly increases the computational cost at inference time, which can be particularly problematic for encoding longer documents. Our approach outperforms Springer et al. (2024), without inducing any additional computational overhead at inference time.
@@ 7 Conclusion
We present LLM2Vec, a strong unsupervised approach to transform any decoder-only LLMs into a (universal) text embedder. We perform an extensive evaluation on word- and sequence-level tasks and demonstrate the effectiveness of LLM2Vec in both unsupervised and supervised settings. Applying LLM2Vec to Mistral-7B achieves a new state-of-the-art performance on MTEB among unsupervised approaches. When combining LLM2Vec with supervised contrastive fine-tuning, Meta-LLaMA-3-8B achieves SOTA performance among approaches that train only on publicly available data (as of May 24, 2024). Beyond our strong empirical contributions, we provide an extensive analysis of how LLM2Vec impacts the underlying model and reveal an intriguing property of Mistral-7B, which explains its strong out of the box performance with bidirectional attention. The simplicity of our approach, as well as its compute and sample-efficiency, makes LLM2vec a promising solution for low-resource and compute constrained scenarios and opens up several interesting avenues for future work.
@@ Appendix A Limitations
The advantages of small bidirectional encoders come at the cost of complex and computationally intensive training regimes (Li et al., 2023a; Xiao et al., 2023; Li et al., 2023a). In contrast, decoder-only models are much more sample-efficient and do not require large-scale contrastive pre-training (Wang et al., 2023). Moreover, the instruction following capabilities of decoder-only models make them strong contenders for building text embedding models that generalize to a wide range of tasks and domains without the need for expensive adaptation.
@@ D.1.1 MNTP training details
For our models, we perform a hyperparameter search to select the percentage of the masked tokens in a sequence choosing from $20\%,40\%,60\%,80\%$, and $90\%$. For each model, we take the best setup (i.e., masking probability and BERT vs. RoBERTa approach) based on the performance on SICK-R (Agirre et al., 2014) task from the MTEB dataset. This results in the following choices: for S-LLaMA-1.3B, LLaMA-2-7B, and Meta-LLaMA-3-8B, we apply BERT’s masking strategy with masking probability of $20\%$. For Mistral-7B, we apply RoBERTa’s masking strategy with probability of $80\%$.
We train all the models for 1000 steps with LoRA $r=16$ and $\alpha=32$, and we follow the same training parameters as RoBERTa MNTP training. When training large 7B and 8B models, we apply brain floating point (bfloat16) quantization, as well as flash attention 2 and gradient checkpointing.
@@ D.1.2 SimCSE training details
Similar to MNTP, we train all models with LoRA $r=16$ and $\alpha=32$ for 1000 steps. For LLaMA-2-7B, Mistral-7B, and Meta-LLaMA-3-8B, we train with a batch size of 128. For S-LLaMA-1.3B, we use a batch size of 32. Additionally, when training LLaMA-2-7B Mistral-7B, and Meta-LLaMA-3-8B, we apply brain floating point (bfloat16) quantization, flash attention 2, and gradient checkpointing.
@@ D.1.3 Word-level training details
We use data from CoNLL-2003, consisting of roughly 14,000 training, 3,250 validation, and 3,450 test samples (Tjong Kim Sang & De Meulder, 2003). We train the classifier for 1,500 steps with a learning rate of $45e-$ 4 and a batch size of 8. For experiments with Mistral-7B models that have been tuned with MNTP, we use the variant which is trained with BERT’s masking strategy and masking probability of 20% (please see D.1.1 for more details). Although 80% masking helps with the performance in sentence-level tasks, it prevents the model from learning proper token representations essential for word-level tasks.
Since the models we experiment with have sub-token based vocabularies, we calculate the embedding of a word by averaging the representations of all its sub-tokens. For example, for a sentence “$w_{1}\;w_{2}\;w_{3}$” which is tokenized as “$\text{BOS}\;\;t_{11}t_{12}\;\;t_{21}t_{22}t_{23}\;\;t_{31}$”, the representation of $w_{1}$, $w_{2}$, and $w_{3}$ will be computed as
@@ Appendix F More analysis results
$q$: the query $s^{+}$: the positive sample $s^{-}$: the negative sample
She loves to travel in summer, especially to cold destinations, avoiding hot and crowded places. She loves to travel in summer, specifically to chilly locations, steering clear of warm, populous areas. She loves to travel in summer, but prefers to visit hot and bustling tourist spots.
The cat often sits by the window, dreaming of chasing birds and enjoying the warm sunshine. The cat often sits by the window, imagining bird pursuits and basking in the sunlight. The cat often sits by the window, but is too lazy to dream of chasing anything.
He reads books every night, finding solace in fiction and escaping from the stresses of daily life. He reads books every night, seeking comfort in stories and evading everyday tensions. He reads books every night, yet he feels that non-fiction is more engaging and informative.
She paints landscapes on weekends, expressing her creativity through vibrant colors and abstract forms. She paints landscapes on weekends, showcasing her artistic flair with lively hues and unconventional shapes. She paints landscapes on weekends, preferring realistic and detailed depictions of nature.
@@
Figure 9 provide cosine similarity results for all three models. In addition to our LLM2Vec-transformed models, we also provide results for Echo emebddings.
@@ F.1 Additional plots
Figure 10 provides additional plots for the representation analysis using two different Wikipedia paragraphs. The trends closely follow those reported in Section 4. Figure 11 shows that the same behavior we observe for Mistral-7B-instruct-v0.2 also holds true for other variants of the Mistral-7B model. We take this as additional evidence that the Mistral-7B base model was trained with some for of bidirectional attention.
@@ G.1 E5 dataset
NLI Given a premise, retrieve a hypothesis that is entailed by the premise
@@
DuReader Given a Chinese search query, retrieve web passages that answer the question
ELI5 Provided a user question, retrieve the highest voted answers on Reddit ELI5 forum
FEVER Given a claim, retrieve documents that support or refute the claim
HotpotQA Given a multi-hop question, retrieve documents that can help answer the question
MIRACL Given a question, retrieve Wikipedia passages that answer the question
MrTyDi Given a question, retrieve Wikipedia passages that answer the question
MSMARCO Passage Given a web search query, retrieve relevant passages that answer the query
MSMARCO Document Given a web search query, retrieve relevant documents that answer the query
NQ Given a question, retrieve Wikipedia passages that answer the question
QuoraDuplicates Given a question, retrieve questions that are semantically equivalent to the given question
@@
SQuAD Retrieve Wikipedia passages that answer the question
T2Ranking Given a Chinese search query, retrieve web passages that answer the question
TriviaQA Retrieve Wikipedia passages that answer the question
@@ G.2 Training details
All models are trained with LoRA $r=16$ and $\alpha=32$, brain floating point (bfloat16) quantization, gradient checkpointing, and flash attention 2 (Dao, 2024) to optimize GPU memory consumption. We train on 8 NVIDIA A100 GPUs with an effective batch size of 512 for 1000 steps using a maximum sequence length of 512 tokens. We use the Adam optimizer with a learning rate of $42e-$ 4 and a linear learning rate warm-up for the first 300 steps.